# LLMs4OM: Matching Ontologies with Large Language Models

### v1

@@ 1 Introduction
This study’s empirical tests are varied not only in their approach but also in the range of ontological knowledge domains they cover. We evaluate LLMs across six tracks of the OAEI campaign, encompassing 20 datasets in total. The primary contributions of this paper are threefold: 1) Introduction of the LLMs4OM, an end-to-end framework that utilizes LLMs for OM; 2) A thorough empirical evaluation of seven state-of-the-art domain-independent LLMs and four retrieval models for their suitability to the various OM tasks; and 3) The source code implementation of the LLMs4OM framework released here https://anonymous.4open.science/r/LLMs4OM.
@@ 6 Conclusion
Supplemental Material Statement:
Our detailed results and codebase are available https://anonymous.4open.science/r/LLMs4OM.

### v2

@@ 1 Introduction
This study’s empirical tests are varied not only in their approach but also in the range of ontological knowledge domains they cover. We evaluate LLMs across six tracks of the OAEI campaign, encompassing 20 datasets in total. The primary contributions of this paper are threefold: 1) Introduction of the LLMs4OM, an end-to-end framework that utilizes LLMs for OM; 2) A thorough empirical evaluation of seven state-of-the-art domain-independent LLMs and four retrieval models for their suitability to the various OM tasks; and 3) The source code implementation of the LLMs4OM framework released here https://github.com/HamedBabaei/LLMs4OM.